a grocer mixes a premium blend worth $17 per kilogram with a blend worth $7 per kilogram to make 36 kilogram blend worth $11 per kilogram. how many kilograms of each type are included?
Answer by jorel1380(3719)

36x11=396
17m+7(36-m)=396
17m+252-7m=396
10m=144
m=14.4
36-14.4=21.6
The grocer needs 14.4 kilos of premium blend and 21.6 kilos of the other blend for his mix..